Let a be a sequence of real numbers and let c ∈ R be a cluster point of a.

Let π:N → N be defined by

π(1) = min{k∈N||ak −c|<1},

π(n+1) = min{k∈N|k>π(n), |ak −c|< 1 } foralln∈N. n+1

(i) Justify the definition of π. (i.e Show that π is well defined.) (ii) Show that π is strictly increasing.

(iii) Prove that the subsequence (aπ(n))N of a converges to c


1
Expert's answer
2021-05-04T13:14:59-0400

Solution :-


Let an be the sequence of real numbers

And c "\epsilon" R

"\Pi : N\rightarrow N"

"\Pi(1) = min{[K\epsilon N || ak-c I < 1]}" ,

"\Pi" (n+1) = min[k∈N|k>"\Pi" (n), |ak −c|< 1 ] for

all n∈N. n+1



(i) as we can see "\Pi" is defined ": N\rightarrow N"


and minimum value of function "\Pi" comes in "\Pi(1) = min{[K\epsilon N || ak-c I < 1]}" ,

and also we can see that "\Pi" (n+1) = min[k∈N|k>"\Pi" (n), |ak −c|< 1 ] for

all n∈N. n+1

we can say that "\Pi" is well defined in the given interval


(ii) we can say that , when we do differentiation of the an and "\Pi" is well defined in the given interval

interval is given for "\Pi"

"\Pi : N\rightarrow N"

"\Pi(1) = min{[K\epsilon N || ak-c I < 1]}" ,

"\Pi" (n+1) = min[k∈N|k>"\Pi" (n), |ak −c|< 1 ] for

all n∈N. n+1

we can say it is differentiaval in the given interval , by differentiation function Π is strictly increasing function


(iii)

 (a"\Pi" (n))N of a converges to c

as we have limits and we can see that

"\Pi : N\rightarrow N"

"\Pi(1) = min{[K\epsilon N || ak-c I < 1]}" ,

"\Pi" (n+1) = min[k∈N|k>"\Pi" (n), |ak −c|< 1 ] for

all n∈N. n+1

by this we can see that every bounded function is convrgence function

{an}n∈N is convergent. Then by Theorem we know that the limit is unique and so we can write it as l

"lim_{n\rightarrow \infin}a_n=l"

or "a_n \rightarrow l"

as "n \rightarrow \infin"

by this we have proved that  the subsequence (a"\Pi" (n))N of a converges to c , AS "\boxed{c \space\epsilon R}"